Solve by elimination:
2/3x + 3/4y = 11/12,
x/3 + 7/18y = 1/2

Answer by waynest(281)

(2/3)x + (3/4)y = (11/12)
(x/3) + (7/18)y = (1/2)
multiply the second equation by 2 and then subtract it from the first
(2/3)x + (3/4)y = (11/12)
(2/3)x + (7/9)y = 1
you will get
-(1/36)y = -(1/12)
-36(-1/36)y = -36(-1/12)
y = 3
substitute 3 for y in the first equation to get x
(2/3)x + (3/4)(3) = (11/12)
(2/3)x + (9/4) = (11/12)
(2/3)x + (9/4)- (9/4) = (11/12) - (9/4)
(2/3)x = (-4/3)
(3/2)(2/3)x = (-4/3)(3/2)
x = -2
check:
(2/3)(-2) + (3/4)(3) = (11/12)
(-4/3) + (9/4) = (11/12)
(11/12) = (11/12)
(-2/3) + (7/18)(3) = (1/2)
(-2/3) + (7/6) = (1/2)
(1/2) = (1/2)